# Chain-of-Thought Prompting Obscures Hallucination Cues in Large Language Models: An Empirical Evaluation

### v1

@@ Appendix G Results
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Textual HaluEval base 65.77 65.67 60.75 69.58 62.34 75.37 53.07 58.65
@@
Semantic HaluEval base 59.76 59.62 62.44 58.34 55.68 65.08 62.86 57.01
@@
Textual PopQA base 75.35 75.38 70.76 65.25 62.90 76.79 66.46 57.58
@@
Semantic PopQA base 71.90 72.13 72.16 53.05 51.68 80.03 72.07 66.24
@@
Textual TriviaQA base 75.56 75.30 70.39 77.20 82.21 64.39 59.43 55.13
@@
Semantic TriviaQA base 79.25 79.22 75.27 59.38 65.61 58.84 68.50 67.35
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Textual HaluEval base 65.90 65.89 63.45 76.71 67.24 57.13 56.49 58.96
@@
Semantic HaluEval base 63.41 63.41 63.49 59.39 58.04 63.39 57.88 60.93
@@
Textual PopQA base 68.01 68.01 62.75 77.25 59.58 64.79 54.91 57.64
@@
Semantic PopQA base 59.55 59.56 56.97 65.23 60.73 59.62 58.84 59.87
@@
Textual TriviaQA base 64.17 64.20 60.37 84.50 73.47 59.00 55.29 57.75
@@
Semantic TriviaQA base 63.57 63.80 59.24 58.61 61.16 56.49 55.69 57.84
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
Textual HaluEval base 0.5287 0.5281 0.6097 0.6464 0.6543 0.6094 0.6346
CoT 0.5142 0.5161 0.5192 0.6450 0.5932 0.5492 0.6156
LtM 0.5063 0.5089 0.5202 0.6276 0.5895 0.5117 0.6502
MRPP 0.5135 0.5129 0.5274 0.6313 0.5773 0.5338 0.6258
Semantic HaluEval base 0.6648 0.6643 0.6543 0.5033 0.5207 0.5382 0.7279
CoT 0.6546 0.6541 0.5513 0.5003 0.5493 0.5061 0.5058
LtM 0.6571 0.6563 0.5830 0.5491 0.5434 0.5019 0.5136
MRPP 0.6295 0.6323 0.5610 0.5389 0.5472 0.5073 0.5079
Textual PopQA base 0.5297 0.5288 0.5387 0.7073 0.6331 0.5134 0.5195
CoT 0.5096 0.5066 0.5372 0.6752 0.5874 0.5386 0.6479
LtM 0.5045 0.5051 0.5504 0.5869 0.5302 0.5647 0.5462
MRPP 0.5015 0.5055 0.5290 0.5659 0.5245 0.5507 0.5403
Semantic PopQA base 0.6584 0.6587 0.6052 0.5721 0.5124 0.5107 0.5606
CoT 0.6410 0.6380 0.6039 0.5019 0.5988 0.6202 0.5179
LtM 0.6010 0.5928 0.5547 0.5189 0.5768 0.5733 0.5319
MRPP 0.5770 0.5744 0.5457 0.5079 0.5514 0.5598 0.5300
Textual TriviaQA base 0.5417 0.5412 0.5716 0.7345 0.7132 0.5357 0.5315
CoT 0.5142 0.5125 0.5257 0.6255 0.5414 0.5362 0.6504
LtM 0.5179 0.5083 0.5178 0.6134 0.5565 0.5152 0.6684
MRPP 0.5244 0.5218 0.5150 0.6263 0.5431 0.5167 0.6592
Semantic TriviaQA base 0.5159 0.5155 0.5261 0.5602 0.6022 0.5857 0.5160
CoT 0.6588 0.6530 0.6027 0.5163 0.5337 0.5707 0.5352
LtM 0.6946 0.6915 0.6295 0.5175 0.5822 0.5937 0.5891
MRPP 0.7004 0.7002 0.6277 0.5286 0.5832 0.6171 0.6046
Textual CNN/Daily Mail base 0.6709 0.6697 0.5414 0.5077 0.5085 0.6315 0.6648
CoT 0.6133 0.6082 0.5757 0.5123 0.5035 0.5274 0.6043
LtM 0.6151 0.6079 0.5758 0.5002 0.5274 0.5576 0.6380
MRPP 0.6145 0.6082 0.5555 0.5068 0.5063 0.5428 0.6256
Semantic CNN/Daily Mail base 0.6769 0.6751 0.5557 0.5258 0.5210 0.5002 0.6716
CoT 0.5823 0.5734 0.5587 0.5075 0.5298 0.5143 0.6942
LtM 0.5919 0.5795 0.5610 0.5027 0.5457 0.5231 0.6877
MRPP 0.6004 0.5901 0.5798 0.5001 0.5242 0.5253 0.6998
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Textual HaluEval base 0.7561 0.7561 0.7311 0.6877 0.6923 0.6182 0.7303 0.5721
@@
Semantic HaluEval base 0.7241 0.7234 0.6998 0.5782 0.5997 0.5938 0.6019 0.5165
@@
Textual PopQA base 0.7127 0.7126 0.6720 0.7293 0.6496 0.6773 0.6106 0.5185
@@
Semantic PopQA base 0.6497 0.6456 0.6578 0.5798 0.5916 0.5955 0.6694 0.5964
@@
Textual TriviaQA base 0.7225 0.7245 0.7337 0.6822 0.7610 0.7329 0.6212 0.6005
@@
Semantic TriviaQA base 0.7305 0.7305 0.7424 0.5611 0.6239 0.5866 0.5993 0.6542
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
@@
Textual base 0.5891 0.5861 0.6295 0.5351 0.5877 0.5139 0.6332
CoT 0.5789 0.5791 0.6059 0.5001 0.5701 0.5386 0.6049
LtM 0.6267 0.6240 0.6444 0.5188 0.5376 0.5425 0.6138
MRPP 0.6049 0.5982 0.6228 0.5033 0.5517 0.5335 0.6040
Semantic
base 0.5681 0.5704 0.6395 0.5260 0.5610 0.5199 0.6695
CoT 0.5531 0.5510 0.5957 0.5278 0.5258 0.5144 0.6625
LtM 0.5873 0.5777 0.6421 0.5521 0.5172 0.5374 0.6677
MRPP 0.5559 0.5527 0.5858 0.5218 0.5089 0.5576 0.6499
@@
Textual base 0.6219 0.6219 0.5960 0.5189 0.5314 0.6053 0.6404
CoT 0.5556 0.5556 0.5264 0.5081 0.5150 0.5404 0.6033
LtM 0.5785 0.5785 0.5583 0.5221 0.5294 0.5392 0.6061
MRPP 0.5564 0.5565 0.5216 0.5131 0.5303 0.5244 0.6043
Semantic base 0.6132 0.6132 0.5899 0.5256 0.5254 0.5249 0.7037
CoT 0.5560 0.5560 0.5296 0.5240 0.5708 0.5058 0.7155
LtM 0.5559 0.5559 0.5527 0.5470 0.5458 0.5141 0.6780
MRPP 0.5315 0.5315 0.5409 0.5202 0.5403 0.5243 0.7055
@@
Textual base 0.5554 0.5569 0.5120 0.5612 0.5396 0.5157 0.5569
CoT 0.5203 0.5203 0.5446 0.5537 0.5473 0.5323 0.5221
LtM 0.5076 0.5074 0.5142 0.5451 0.5045 0.5978 0.5221
MRPP 0.5107 0.5111 0.5771 0.5891 0.5590 0.5769 0.5120
Semantic base 0.5588 0.5583 0.5455 0.5125 0.5261 0.6004 0.7052
CoT 0.5293 0.5289 0.5072 0.5177 0.5298 0.5893 0.6716
LtM 0.5216 0.5219 0.5062 0.5056 0.5248 0.5674 0.6806
MRPP 0.5558 0.5559 0.5563 0.5642 0.5649 0.5368 0.6449
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Textual HaluEval base -27.09 -27.08 -22.92 45.51 -29.37 16.40 11.06 15.26
@@
Semantic HaluEval base -21.87 -21.86 -22.16 16.47 -16.01 26.81 15.33 20.13
@@
Textual PopQA base -29.69 -29.70 -21.02 44.93 -15.79 29.29 8.10 12.60
@@
Semantic PopQA base -16.99 -16.99 -13.97 29.24 -22.63 15.92 13.69 19.34
@@
Textual TriviaQA base -19.28 -19.33 -14.11 46.94 -31.94 20.66 7.20 10.55
@@
Semantic TriviaQA base -27.44 -27.77 -17.31 13.55 -19.10 17.32 11.50 17.15
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score
Textual HaluEval base -0.0378 -0.0369 -0.1444 0.1928 -0.2031 0.1440 0.1772
@@
Semantic HaluEval base -0.3112 -0.3104 -0.2987 0.0689 -0.0066 0.0434 0.3913
@@
Textual PopQA base -0.0377 -0.0366 -0.0492 0.2638 -0.1694 0.0171 0.0249
@@
Semantic PopQA base -0.2394 -0.2408 -0.1381 0.1137 -0.0389 0.1260 0.0708
@@
Textual TriviaQA base -0.0722 -0.0714 -0.1241 0.4061 -0.3693 0.0619 0.0546
@@
Semantic TriviaQA base -0.0237 -0.0231 -0.0419 0.0866 -0.2094 0.2345 0.0126
@@
Textual CNN/Daily Mail base -0.3460 -0.3437 -0.0701 0.0140 -0.0183 0.2672 0.3268
@@
Semantic CNN/Daily Mail base -0.3661 -0.3631 -0.1145 0.0553 -0.0312 0.0001 0.3559
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score Attn Score
Textual HaluEval base -0.4222 -0.4222 -0.4078 0.3499 -0.3632 0.2789 0.2739 0.0287
@@
Semantic HaluEval base -0.4651 -0.4638 -0.4184 0.1525 -0.2041 0.2288 0.1875 0.1358
@@
Textual PopQA base -0.3591 -0.3591 -0.2905 0.2583 -0.2526 0.3334 0.1868 0.3130
@@
Semantic PopQA base -0.3100 -0.3014 -0.3241 0.1378 -0.2146 0.1757 0.3143 0.2033
@@
Textual TriviaQA base -0.2103 -0.2121 -0.2208 0.1721 -0.2467 0.2355 0.1145 0.0950
@@
Semantic TriviaQA base -0.3406 -0.3407 -0.4543 0.0085 -0.1633 0.1357 0.1771 0.2031
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
Textual HaluEval base 16.89 17.00 24.90 8.40 23.64 25.90 40.09
@@
Semantic HaluEval base 19.50 19.34 13.37 16.08 21.90 25.24 37.27
@@
Textual PopQA base 20.81 21.18 12.67 9.69 22.75 40.21 53.10
@@
Semantic PopQA base 15.29 15.61 4.98 16.66 27.78 27.39 41.15
@@
Textual TriviaQA base 23.98 24.30 27.40 40.62 22.30 10.65 10.93
@@
Semantic TriviaQA base 11.20 12.07 11.90 24.04 21.31 10.47 20.79
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
Textual HaluEval base 12.80 12.75 23.43 19.43 17.78 18.71 23.51
@@
Semantic HaluEval base 11.26 11.23 19.60 26.37 23.74 14.24 18.15
@@
Textual PopQA base 17.76 17.74 26.49 22.52 28.24 21.89 25.30
@@
Semantic PopQA base 14.19 14.15 18.56 22.49 25.23 15.83 18.74
@@
Textual TriviaQA base 31.75 31.82 13.52 6.04 30.41 25.51 22.26
@@
Semantic TriviaQA base 11.45 11.98 23.09 36.09 25.21 14.18 14.59
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Verbalized Certainty Hidden Score
Textual HaluEval base 0.2133 0.2098 0.4199 0.5159 0.3615 0.3567 0.4295
@@
Semantic HaluEval base 0.2590 0.2607 0.1069 0.1943 0.2234 0.1299 0.1212
@@
Textual PopQA base 0.2064 0.2054 0.4866 0.3780 0.4322 0.3659 0.5111
@@
Semantic PopQA base 0.2273 0.2270 0.2293 0.2563 0.3350 0.2059 0.1713
@@
Textual TriviaQA base 0.1205 0.1216 0.1923 0.2668 0.1416 0.1565 0.2142
@@
Semantic TriviaQA base 0.2113 0.2074 0.2000 0.2514 0.2351 0.1268 0.1734
@@
Textual CNN/Daily Mail base 0.1048 0.1093 0.3753 0.4203 0.3140 0.1259 0.2378
@@
Semantic CNN/Daily Mail base 0.0950 0.0969 0.2595 0.3013 0.4183 0.1774 0.1289
@@
PPL Sharpness Eigen Score SelfCk -Prompt SelfCk -NLI Hidden Score Attn Score
Textual HaluEval base 0.1653 0.1650 0.1446 0.0923 0.1761 0.3062 0.3536
@@
Semantic HaluEval base 0.1075 0.0977 0.1295 0.2224 0.2170 0.1082 0.2587
@@
Textual PopQA base 0.1507 0.1508 0.1232 0.0992 0.1803 0.2715 0.4548
@@
Semantic PopQA base 0.1322 0.1274 0.0769 0.1423 0.2190 0.1820 0.4065
@@
Textual TriviaQA base 0.4652 0.4484 0.4546 0.4893 0.4617 0.2768 0.1262
@@
Semantic TriviaQA base 0.1657 0.1853 0.0907 0.2557 0.2150 0.1129 0.1556

### v3

@@ Appendix G Results
G.1 MIND Evaluation Details
To explore whether our observations extend to model-training-based detection paradigms, we conducted additional experiments using the MIND framework (Su et al., 2024).
MIND is an unsupervised, internal-state-based hallucination detection framework. Its key advantages include:
  • Leveraging pseudo-labeled data automatically constructed from Wikipedia for training, thus avoiding manual annotation.
  • Training lightweight classifiers on the LLM’s hidden states for real-time detection.
However, it also shares limitations with other model-training-based approaches:
  • It requires training an auxiliary detector tailored to each specific target LLM.
  • Its performance is dependent on the design and quality of the pseudo-labeled data.
This need for model-specific training is why we primarily focused on non-training-based detection methods in our main experiments.
Experimental Setup. Due to the tight coupling between MIND’s released code and its original LLM architecture, our evaluation was limited to the LLaMA-2-7b-hf model. We evaluated its performance on the HaluEval dataset under different prompting scenarios: standard prompting (base), and the three CoT prompting methods studied in this work.
Results. The results, presented in Table 5, show that MIND’s detection performance also degrades when the LLM responds after CoT prompting. This degradation is consistent across both Textual and Semantic hallucination annotations and across both AUROC and Correlation metrics, reinforcing the broad and challenging nature of the phenomenon we identify.
Table 5: Performance of the MIND detector on HaluEval under different prompting scenarios.
Textual Semantic
AUROC Corr AUROC Corr
base 0.6765 0.4733 0.5904 0.2941
CoT 0.6521 0.4261 0.5572 0.2307
MRPP 0.6513 0.4304 0.5608 0.2381
LtM 0.6379 0.3983 0.5573 0.2246
G.2 Detail of Results
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Verbalized
Certainty
Hidden
Score
Attn
Score
Textual
HaluEval
base 65.77 65.67 60.75 69.58 62.34 75.37 53.07 58.65
@@
Semantic
HaluEval
base 59.76 59.62 62.44 58.34 55.68 65.08 62.86 57.01
@@
Textual
PopQA
base 75.35 75.38 70.76 65.25 62.90 76.79 66.46 57.58
@@
Semantic
PopQA
base 71.90 72.13 72.16 53.05 51.68 80.03 72.07 66.24
@@
Textual
TriviaQA
base 75.56 75.30 70.39 77.20 82.21 64.39 59.43 55.13
@@
Semantic
TriviaQA
base 79.25 79.22 75.27 59.38 65.61 58.84 68.50 67.35
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Verbalized
Certainty
Hidden
Score
Attn
Score
Textual
HaluEval
base 65.90 65.89 63.45 76.71 67.24 57.13 56.49 58.96
@@
Semantic
HaluEval
base 63.41 63.41 63.49 59.39 58.04 63.39 57.88 60.93
@@
Textual
PopQA
base 68.01 68.01 62.75 77.25 59.58 64.79 54.91 57.64
@@
Semantic
PopQA
base 59.55 59.56 56.97 65.23 60.73 59.62 58.84 59.87
@@
Textual
TriviaQA
base 64.17 64.20 60.37 84.50 73.47 59.00 55.29 57.75
@@
Semantic
TriviaQA
base 63.57 63.80 59.24 58.61 61.16 56.49 55.69 57.84
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Hidden
Score
Verbalized
Certainty
Attn
Score
Textual
HaluEval
base 0.5287 0.5281 0.6097 0.6464 0.6543 0.6094 0.6094 0.6346
CoT 0.5142 0.5161 0.5192 0.6450 0.5932 0.5492 0.5492 0.6156
LtM 0.5063 0.5089 0.5202 0.6276 0.5895 0.5117 0.5117 0.6502
MRPP 0.5135 0.5129 0.5274 0.6313 0.5773 0.5338 0.5338 0.6258
Semantic
HaluEval
base 0.6648 0.6643 0.6543 0.5033 0.5207 0.5382 0.5382 0.7279
CoT 0.6546 0.6541 0.5513 0.5003 0.5493 0.5061 0.5061 0.5058
LtM 0.6571 0.6563 0.5830 0.5491 0.5434 0.5019 0.5019 0.5136
MRPP 0.6295 0.6323 0.5610 0.5389 0.5472 0.5073 0.5073 0.5079
Textual
PopQA
base 0.5297 0.5288 0.5387 0.7073 0.6331 0.5134 0.5134 0.5195
CoT 0.5096 0.5066 0.5372 0.6752 0.5874 0.5386 0.5386 0.6479
LtM 0.5045 0.5051 0.5504 0.5869 0.5302 0.5647 0.5647 0.5462
MRPP 0.5015 0.5055 0.5290 0.5659 0.5245 0.5507 0.5507 0.5403
Semantic
PopQA
base 0.6584 0.6587 0.6052 0.5721 0.5124 0.5107 0.5107 0.5606
CoT 0.6410 0.6380 0.6039 0.5019 0.5988 0.6202 0.6202 0.5179
LtM 0.6010 0.5928 0.5547 0.5189 0.5768 0.5733 0.5733 0.5319
MRPP 0.5770 0.5744 0.5457 0.5079 0.5514 0.5598 0.5598 0.5300
Textual
TriviaQA
base 0.5417 0.5412 0.5716 0.7345 0.7132 0.5357 0.5357 0.5315
CoT 0.5142 0.5125 0.5257 0.6255 0.5414 0.5362 0.5362 0.6504
LtM 0.5179 0.5083 0.5178 0.6134 0.5565 0.5152 0.5152 0.6684
MRPP 0.5244 0.5218 0.5150 0.6263 0.5431 0.5167 0.5167 0.6592
Semantic
TriviaQA
base 0.5159 0.5155 0.5261 0.5602 0.6022 0.5857 0.5857 0.5160
CoT 0.6588 0.6530 0.6027 0.5163 0.5337 0.5707 0.5707 0.5352
LtM 0.6946 0.6915 0.6295 0.5175 0.5822 0.5937 0.5937 0.5891
MRPP 0.7004 0.7002 0.6277 0.5286 0.5832 0.6171 0.6171 0.6046
Textual
CNN/Daily Mail
base 0.6709 0.6697 0.5414 0.5077 0.5085 0.6315 0.6315 0.6648
CoT 0.6133 0.6082 0.5757 0.5123 0.5035 0.5274 0.5274 0.6043
LtM 0.6151 0.6079 0.5758 0.5002 0.5274 0.5576 0.5576 0.6380
MRPP 0.6145 0.6082 0.5555 0.5068 0.5063 0.5428 0.5428 0.6256
Semantic
CNN/Daily Mail
base 0.6769 0.6751 0.5557 0.5258 0.5210 0.5002 0.5002 0.6716
CoT 0.5823 0.5734 0.5587 0.5075 0.5298 0.5143 0.5143 0.6942
LtM 0.5919 0.5795 0.5610 0.5027 0.5457 0.5231 0.5231 0.6877
MRPP 0.6004 0.5901 0.5798 0.5001 0.5242 0.5253 0.5253 0.6998
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Verbalized
Certainty
Hidden
Score
Attn
Score
Textual
HaluEval
base 0.7561 0.7561 0.7311 0.6877 0.6923 0.6182 0.7303 0.5721
@@
Semantic
HaluEval
base 0.7241 0.7234 0.6998 0.5782 0.5997 0.5938 0.6019 0.5165
@@
Textual
PopQA
base 0.7127 0.7126 0.6720 0.7293 0.6496 0.6773 0.6106 0.5185
@@
Semantic
PopQA
base 0.6497 0.6456 0.6578 0.5798 0.5916 0.5955 0.6694 0.5964
@@
Textual
TriviaQA
base 0.7225 0.7245 0.7337 0.6822 0.7610 0.7329 0.6212 0.6005
@@
Semantic
TriviaQA
base 0.7305 0.7305 0.7424 0.5611 0.6239 0.5866 0.5993 0.6542
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Hidden
Score
Verbalized
Certainty
Attn
Score
@@
Textual base 0.5891 0.5861 0.6295 0.5351 0.5877 0.5139 0.5398 0.6332
CoT 0.5789 0.5791 0.6059 0.5001 0.5701 0.5386 0.5310 0.6049
LtM 0.6267 0.6240 0.6444 0.5188 0.5376 0.5425 0.5354 0.6138
MRPP 0.6049 0.5982 0.6228 0.5033 0.5517 0.5335 0.5409 0.6040
Semantic base 0.5681 0.5704 0.6395 0.5260 0.5610 0.5199 0.5287 0.6695
CoT 0.5531 0.5510 0.5957 0.5278 0.5258 0.5144 0.5320 0.6625
LtM 0.5873 0.5777 0.6421 0.5521 0.5172 0.5374 0.5085 0.6677
MRPP 0.5559 0.5527 0.5858 0.5218 0.5089 0.5576 0.5245 0.6499
@@
Textual base 0.6219 0.6219 0.5960 0.5189 0.5314 0.6053 0.5487 0.6404
CoT 0.5556 0.5556 0.5264 0.5081 0.5150 0.5404 0.5362 0.6033
LtM 0.5785 0.5785 0.5583 0.5221 0.5294 0.5392 0.5344 0.6061
MRPP 0.5564 0.5565 0.5216 0.5131 0.5303 0.5244 0.5410 0.6043
Semantic base 0.6132 0.6132 0.5899 0.5256 0.5254 0.5249 0.5316 0.7037
CoT 0.5560 0.5560 0.5296 0.5240 0.5708 0.5058 0.5420 0.7155
LtM 0.5559 0.5559 0.5527 0.5470 0.5458 0.5141 0.5312 0.6780
MRPP 0.5315 0.5315 0.5409 0.5202 0.5403 0.5243 0.5302 0.7055
@@
Textual base 0.5554 0.5569 0.5120 0.5612 0.5396 0.5157 0.5313 0.5569
CoT 0.5203 0.5203 0.5446 0.5537 0.5473 0.5323 0.5249 0.5221
LtM 0.5076 0.5074 0.5142 0.5451 0.5045 0.5978 0.5299 0.5221
MRPP 0.5107 0.5111 0.5771 0.5891 0.5590 0.5769 0.5421 0.5120
Semantic base 0.5588 0.5583 0.5455 0.5125 0.5261 0.6004 0.5308 0.7052
CoT 0.5293 0.5289 0.5072 0.5177 0.5298 0.5893 0.5238 0.6716
LtM 0.5216 0.5219 0.5062 0.5056 0.5248 0.5674 0.5108 0.6806
MRPP 0.5558 0.5559 0.5563 0.5642 0.5649 0.5368 0.5240 0.6449
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Verbalized
Certainty
Hidden
Score
Attn
Score
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Verbalized
Certainty
Hidden
Score
Attn
Score
Textual
HaluEval
base -27.09 -27.08 -22.92 45.51 -29.37 16.40 11.06 15.26
@@
Semantic
HaluEval
base -21.87 -21.86 -22.16 16.47 -16.01 26.81 15.33 20.13
@@
Textual
PopQA
base -29.69 -29.70 -21.02 44.93 -15.79 29.29 8.10 12.60
@@
Semantic
PopQA
base -16.99 -16.99 -13.97 29.24 -22.63 15.92 13.69 19.34
@@
Textual
TriviaQA
base -19.28 -19.33 -14.11 46.94 -31.94 20.66 7.20 10.55
@@
Semantic
TriviaQA
base -27.44 -27.77 -17.31 13.55 -19.10 17.32 11.50 17.15
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Verbalized
Certainty
Hidden
Score
Textual
HaluEval
base -0.0378 -0.0369 -0.1444 0.1928 -0.2031 0.1440 0.1772
@@
Semantic
HaluEval
base -0.3112 -0.3104 -0.2987 0.0689 -0.0066 0.0434 0.3913
@@
Textual
PopQA
base -0.0377 -0.0366 -0.0492 0.2638 -0.1694 0.0171 0.0249
@@
Semantic
PopQA
base -0.2394 -0.2408 -0.1381 0.1137 -0.0389 0.1260 0.0708
@@
Textual
TriviaQA
base -0.0722 -0.0714 -0.1241 0.4061 -0.3693 0.0619 0.0546
@@
Semantic
TriviaQA
base -0.0237 -0.0231 -0.0419 0.0866 -0.2094 0.2345 0.0126
@@
Textual
CNN/Daily Mail
base -0.3460 -0.3437 -0.0701 0.0140 -0.0183 0.2672 0.3268
@@
Semantic
CNN/Daily Mail
base -0.3661 -0.3631 -0.1145 0.0553 -0.0312 0.0001 0.3559
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Verbalized
Certainty
Hidden
Score
Attn
Score
Textual
HaluEval
base -0.4222 -0.4222 -0.4078 0.3499 -0.3632 0.2789 0.2739 0.0287
@@
Semantic
HaluEval
base -0.4651 -0.4638 -0.4184 0.1525 -0.2041 0.2288 0.1875 0.1358
@@
Textual
PopQA
base -0.3591 -0.3591 -0.2905 0.2583 -0.2526 0.3334 0.1868 0.3130
@@
Semantic
PopQA
base -0.3100 -0.3014 -0.3241 0.1378 -0.2146 0.1757 0.3143 0.2033
@@
Textual
TriviaQA
base -0.2103 -0.2121 -0.2208 0.1721 -0.2467 0.2355 0.1145 0.0950
@@
Semantic
TriviaQA
base -0.3406 -0.3407 -0.4543 0.0085 -0.1633 0.1357 0.1771 0.2031
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Hidden
Score
Attn
Score
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Hidden
Score
Attn
Score
Textual
HaluEval
base 16.89 17.00 24.90 8.40 23.64 25.90 40.09
@@
Semantic
HaluEval
base 19.50 19.34 13.37 16.08 21.90 25.24 37.27
@@
Textual
PopQA
base 20.81 21.18 12.67 9.69 22.75 40.21 53.10
@@
Semantic
PopQA
base 15.29 15.61 4.98 16.66 27.78 27.39 41.15
@@
Textual
TriviaQA
base 23.98 24.30 27.40 40.62 22.30 10.65 10.93
@@
Semantic
TriviaQA
base 11.20 12.07 11.90 24.04 21.31 10.47 20.79
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Hidden
Score
Attn
Score
Textual
HaluEval
base 12.80 12.75 23.43 19.43 17.78 18.71 23.51
@@
Semantic
HaluEval
base 11.26 11.23 19.60 26.37 23.74 14.24 18.15
@@
Textual
PopQA
base 17.76 17.74 26.49 22.52 28.24 21.89 25.30
@@
Semantic
PopQA
base 14.19 14.15 18.56 22.49 25.23 15.83 18.74
@@
Textual
TriviaQA
base 31.75 31.82 13.52 6.04 30.41 25.51 22.26
@@
Semantic
TriviaQA
base 11.45 11.98 23.09 36.09 25.21 14.18 14.59
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Verbalized
Certainty
Hidden
Score
Textual
HaluEval
base 0.2133 0.2098 0.4199 0.5159 0.3615 0.3567 0.4295
@@
Semantic
HaluEval
base 0.2590 0.2607 0.1069 0.1943 0.2234 0.1299 0.1212
@@
Textual
PopQA
base 0.2064 0.2054 0.4866 0.3780 0.4322 0.3659 0.5111
@@
Semantic
PopQA
base 0.2273 0.2270 0.2293 0.2563 0.3350 0.2059 0.1713
@@
Textual
TriviaQA
base 0.1205 0.1216 0.1923 0.2668 0.1416 0.1565 0.2142
@@
Semantic
TriviaQA
base 0.2113 0.2074 0.2000 0.2514 0.2351 0.1268 0.1734
@@
Textual
CNN/Daily Mail
base 0.1048 0.1093 0.3753 0.4203 0.3140 0.1259 0.2378
@@
Semantic
CNN/Daily Mail
base 0.0950 0.0969 0.2595 0.3013 0.4183 0.1774 0.1289
@@
PPL Sharpness
Eigen
Score
SelfCk
-Prompt
SelfCk
-NLI
Hidden
Score
Attn
Score
Textual
HaluEval
base 0.1653 0.1650 0.1446 0.0923 0.1761 0.3062 0.3536
@@
Semantic
HaluEval
base 0.1075 0.0977 0.1295 0.2224 0.2170 0.1082 0.2587
@@
Textual
PopQA
base 0.1507 0.1508 0.1232 0.0992 0.1803 0.2715 0.4548
@@
Semantic
PopQA
base 0.1322 0.1274 0.0769 0.1423 0.2190 0.1820 0.4065
@@
Textual
TriviaQA
base 0.4652 0.4484 0.4546 0.4893 0.4617 0.2768 0.1262
@@
Semantic
TriviaQA
base 0.1657 0.1853 0.0907 0.2557 0.2150 0.1129 0.1556